Circular No. 9118

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A. cbatiau@eps.harvard.edu ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET P/2010 D1 (WISE) A. Mainzer, Jet Propulsion Laboratory, reports that an object discov-

ered on 12- and 22-μm images by the WISE satellite (discovery observation tabulated below) shows a ‘hazy’ coma at least 35′′ in diameter with a hint of a faint tail; it shows a less-pronounced nuclear condensation than the WISE images of P/2010 B2 (cf. IAUC 9115) and P/2009 WJ_{50} (cf. IAUC 9117). R. S. McMillan, University of Arizona, reports that Spacewatch CCD images taken with the 1.8-m f/2.7 reflector on Feb. 19.2 UT show the object to be more diffuse than stars, despite poor seeing, of magnitude V = 21.1–21.5, with an extension 10′′–15′′ long toward p.a. 80^{o}.

2010 UT  α_{2000}  δ_{2000}  Observer
Feb. 17.34195 4^{h}19^{m}12.^{s}06 +10^{o}55′42′′.7 WISE

The available astrometry (including prediscovery Catalina Sky Survey observations from 2009 Nov. 9 and Dec. 10), the following elliptical orbital elements, residuals, and an ephemeris appear on MPEC 2010-D51.

Epoch = 2009 June 18.0 TT

T = 2009 June 25.5899 TT  ω = 225.^{o}9003
e = 0.356553  Ω = 160.9008 }2000.0
q = 2.668963 AU  i = 9.6471

a = 4.147912 AU n^{o} = 0.1166702 P = 8.45 years

V496 SCUTI AND V5584 SAGITTARII R. W. Russell, E. A. Laag, and R. J. Rudy, Aerospace Corporation;

and M. A. Skinner and S. A. Gregory, Boeing Corporation, report 3- to 14-μm spectroscopy of V496 Scuti (cf. IAUC 9093) and V5584 Sgr (cf. IAUC 9089) on Feb. 10 using the Aerospace Corporation’s Broadband Array Spectrograph System on the 3.6-m AEOS telescope. Both novae have formed dust since previous observations in December. Thermal emission from this dust dominates the infrared continuum in both objects. The dust temperatures are 1200 ± 100 K for V496 Sct and 880 ± 50 K for V5584 Sgr. There is no evidence for carbon-monoxide emission in V496 Sct, and it is likely that the period of CO emission reported previously for this object (IAUC 9099, CBET 2069) has ended.

2010 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green